# Bidirectional Nonlinear Optical Tomography: Unbiased Characterization of Off- and On-Chip Coupling Efficiencies

### v1

@@ Abstract
Accurate evaluation of nonlinear photonic integrated circuits requires separating input and output coupling efficiencies (i.e., $\eta_{1}$ and $\eta_{2}$), yet the conventional linear-transmission calibration method recovers only their product (i.e., $\eta_{1}\eta_{2}$) and therefore introduces systematic bias when inferring on-chip performance from off-chip data. We present bidirectional nonlinear optical tomography (BNOT), a direction-aware metrology that uses forward and backward pumping of complementary nonlinear probes, with process-appropriate detection, to break the “degeneracy” of $\eta_{1}\eta_{2}$ and estimate individual interface efficiencies with tight confidence intervals. The method links off-chip measurements to on-chip quantities through a compact observation model that explicitly incorporates pump fluctuations and detector noise, and it frames efficiency extraction as a joint constrained optimization. Monte Carlo studies show unbiased convergence of the estimated efficiencies to ground truth with low error across realistic operating regimes. Using these efficiency estimates to reconstruct on-chip nonlinear figures of merit yields distributions centered on the true values with reduced variance, whereas conventional “degenerate” calibration is biased and can substantially misestimate on-chip performance. BNOT is hardware-compatible and platform-agnostic, and it provides unbiased unbiased characterization of off- and on-chip coupling efficiencies across nonlinear processes, enabling reproducible, coupling-resolved benchmarking for scalable systems in quantum optics, frequency conversion, and precision metrology.
@@
Nonlinear photonic integrated circuits (PICs) promise a compact wafer-scale platform that transforms passive chips into active nonlinear engines for frequency conversion and generation of quantum states [1, 2]. Across emerging platforms such as lithium niobate (LiN), silicon nitride (SiN), aluminum nitride, and gallium arsenide, canonical processes, including spontaneous parametric down conversions (SPDC) [3, 4, 5], four-wave mixing (FWM) [6, 7, 8, 9, 10, 11, 12, 13], harmonic generation [14, 15, 16, 17], self-phase modulation (SPM) [18], optical parametric oscillation (OPO) [19, 20, 21, 22] and optical parametric amplification (OPA) [23, 3, 16], are now accessible on chip, enabling applications in spectroscopy, sensing, and quantum information science [3, 6, 7, 8, 9, 14, 15, 19, 20, 24]. These processes are carried out in diverse PIC structures (Fig. 1(a)), ranging from waveguides to a variety of microring resonators, which provide tailored dispersion and mode confinement. Multimode and microcomb architectures further scale brightness and mode count with favorable hardware efficiency [25], positioning nonlinear PICs as strong candidates for continuous-variable (CV) quantum computing [7, 26, 27] and classical optical computing [28, 29, 30, 31].
@@
A major challenge in nonlinear PICs is the lack of bias-free performance estimation, especially as nonlinear interaction strength increases. Because experiments typically rely on off-chip lasers and detectors, the reported metrics involve two coupling interfaces whose efficiencies $(\eta_{1},\eta_{2})$ are often unequal due to mode mismatch, polarization sensitivity, fabrication imperfections, or differing coupling schemes (e.g., edge coupler on one side and grating coupler on the other). Conventional linear-transmission measurements yield only the product $\propto\eta_{1}\eta_{2}$, obscuring which interface dominates the loss and leading to the “degenerate” estimate $(\eta_{1}\eta_{2})^{1/2}$, a source of systematic bias in evaluating on-chip performance [32, 33, 34, 32]. An example is chip-based second harmonic generation (SHG) measured off chip, where the detected efficiency depends asymmetrically on the two interfaces. This can be expressed as $f_{AS}(\eta_{1},\eta_{2})$, satisfying $f_{AS}(\eta_{1},\eta_{2})\neq f_{AS}(\eta_{2},\eta_{1})$ and scaling as $P_{\text{SHG}}/P_{\text{in}}^{2}\propto\eta_{\text{in}}^{2}\eta_{\text{out}}$ [35]. Conventional calibration assumes $\eta_{\text{in}}=\eta_{\text{out}}=(\eta_{1}\eta_{2})^{1/2}$, giving $P_{\text{SHG}}/P_{\text{in}}^{2}\propto(\eta_{1}\eta_{2})^{3/2}$ (curve (i) in Fig. 1(b)), whereas the actual response depends on direction: forward pumping ($\eta_{\text{in}}=\eta_{1}$, $\eta_{\text{out}}=\eta_{2}$) yields $\propto\eta_{1}^{2}\eta_{2}$ (curve (ii)), and backward pumping ($\eta_{\text{in}}=\eta_{2}$, $\eta_{\text{out}}=\eta_{1}$) yields $\propto\eta_{2}^{2}\eta_{1}$ (curve (iii)). This intrinsic asymmetry ultimately explains the systematic deviation observed under symmetric calibration. Wang et al. [36] used forward and backward SHG as symmetric and asymmetric references to attribute performance differences to the different input/output coupling efficiencies. However, such approaches still assume equal efficiencies in the symmetric case and cannot uniquely resolve individual efficiencies.
@@
Among nonlinear processes, squeezed light provides another stringent benchmark for on-chip device performance. Bulk-optics squeezing experiment have reached $15\penalty 10000\ \text{dB}$ [37], while envisioned applications require near-$10\penalty 10000\ \text{dB}$ for CV fault-tolerant quantum computing [26], gravitational-wave detectors [38], and Gottesman-Kitaev-Preskill (GKP) sources [39]. Integrated platforms have advanced from $1.7\penalty 10000\ \text{dB}$ demonstrations [10] to multi-decibel squeezing in LiN [7], cavity optomechanics [40], and foundry-compatible SiN [41, 12], with further gains in nanophotonic molecules and microcombs [42, 11]. The measured squeezing levels reported (off-chip) now exceed $4.9\penalty 10000\ \text{dB}$ in PPLN waveguides, $3.5\penalty 10000\ \text{dB}$ in Kerr microrings, $5.6\penalty 10000\ \text{dB}$ in microcombs [3, 6, 9], and attain $3.1\penalty 10000\ \text{dB}$ in wafer-scale integration [12]. As measured values increase and target on-chip squeezing thresholds approach $10-15\penalty 10000\ \text{dB}$, an accurate estimation of interface efficiencies $(\eta_{1},\eta_{2})$ is becoming even more critical [43, 44, 6, 45].
@@
Figure 2: Histograms of Monte Carlo (MC) simulations with the ground-truth interface efficiencies in Eq. (5) and parameters in Tab. 1. (a) Simulated off-chip squeezing and (b) SHG efficiency. (c) BNOT estimates of $\hat{\eta}_{1}$ (blue) and $\hat{\eta}_{2}$ (red), with shaded regions showing 95 % CIs (2.5–97.5 percentiles).
I Concept
Our BNOT methodology leverages the reversibility of two nonlinear processes within the same photonic device by pumping it in opposite directions. To benchmark its performance, we consider on-chip squeezed-light generation and SHG in an exemplary platform: periodically poled LiN (PPLN) waveguides. The calibration scheme is illustrated in Fig. 1(c). A pump field at $2\omega$ is injected into the PPLN waveguide through the left interface in the forward direction (left-to-right), while a pump at $\omega$ is simultaneously launched through the right interface in the backward direction (right-to-left). The corresponding interface coupling efficiencies are denoted $\eta_{1}^{(2\omega)}$, $\eta_{2}^{(2\omega)}$, $\eta_{1}^{(\omega)}$, and $\eta_{2}^{(\omega)}$, where the superscript denotes the pump frequency. This bidirectional pumping exploits the reversibility of the nonlinear interactions, enabling both degenerate SPDC and SHG to occur concurrently. The corresponding output fields are measured at their respective ports using balanced homodyne detection and direct detection.
To relate on- and off-chip performances in squeezing and SHG, we define two estimation variables, $x_{1}$ and $x_{2}$, corresponding to the left- and right-coupling efficiencies of the PPLN waveguide. Notably, these are not the ground-truth efficiencies: $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ but the tunable parameters in our model used to recover them. The off-chip squeezing level (in dB) and SHG efficiency (in W-1m-2) are expressed as
@@ I Concept
$\mathcal{E}_{\text{OFF}}\,(x_{1},x_{2})$ $=x_{1}\,\mathcal{E}_{\text{ON}}(x_{2})+\xi_{\text{shg}},$
where
$\mathcal{S}_{\text{ON}}\,(x_{1})$ $=2\sqrt{x_{1}\,\alpha\,P_{2\omega}\,\varepsilon_{1}}\,\mathrm{sinc}\left(\Delta k\,L/2\right)\;\;\;\;\text{and}$ (2)
$\mathcal{E}_{\text{ON}}\,(x_{2})$ $=x_{2}^{2}\,\varepsilon_{2}^{2}\,\alpha\,\mathrm{sinc}^{2}\left(\Delta k\,L/2\right)$
denote the analytical formula for the on-chip squeezing [46] and the SHG efficiency [17]. Here $P_{2\omega}$ denotes the non-depleted off-chip pump power at $2\omega$ for forward pumping, $\varepsilon_{1}\sim U(1-\delta_{1},1+\delta_{1})$ and $\varepsilon_{2}\sim U(1-\delta_{2},1+\delta_{2})$ are the uniform fluctuations of the corresponding pump powers, $\delta_{1}$ and $\delta_{2}$ are the half of the corresponding fluctuation ranges, $\xi_{\text{sqz}}\sim\mathcal{N}(0,\sigma^{2}_{\text{sqz}})$ and $\xi_{\text{shg}}\sim\mathcal{N}(0,\sigma^{2}_{\text{shg}})$ represent additive Gaussian noise in the squeezing and SHG measurements (e.g., instability of the measurement system), and $\alpha=8\pi^{2}d_{\mathrm{eff}}^{2}L^{2}/\varepsilon_{0}cn_{2\omega}n_{\omega}^{2}\lambda_{\omega}^{2}A$. The parameter $d_{\mathrm{eff}}$ is the effective nonlinear coefficient under quasi-phase-matching, $L$ is the poling length, $\varepsilon_{0}$ is the vacuum permittivity, $c$ is the speed of light, $n_{2\omega}$ and $n_{\omega}$ are the refractive indices of lithium niobate at $2\omega$ and $\omega$, respectively, and $\Delta k$ and $A$ denote the wave-vector mismatch and waveguide mode area (i.e., $\lambda_{\omega}=2\pi c/\omega$).
@@
Figure 3: (a) Total root-mean-square error (RMSE) map of efficiency estimation using BNOT, $\sqrt{e_{1}^{2}+e_{2}^{2}}$, where $e_{1}$ and $e_{2}$ are the RMSEs of $\eta^{(2\omega)}_{1}$ and $\eta^{(\omega)}_{2}$, respectively. The red dashed contour marks the $0.01$ boundary. Colored stars indicate the parameter sets used for MC simulations. (b) RMSE of $\eta^{(2\omega)}_{1}$ and (c) RMSE of $\eta^{(\omega)}_{2}$ as functions of the number of independent trials, with colors matching the parameter sets in (a).
@@
The conventional calibration approach based on linear transmission cannot distinguish between left-to-right and right-to-left pumping [32, 33, 34, 32]. As a result, it enforces a “degenerate”-interface assumption, estimating the efficiencies as
$\hat{\eta}_{1,\text{C}}=\hat{\eta}_{2,\text{C}}$ $=\left(\eta_{2}^{(\omega)}\eta_{1}^{(\omega)}\right)^{1/2}\;\;\;\;\;\,\text{at $\omega$},$ (3)
$=\left(\eta_{2}^{(2\omega)}\eta_{1}^{(2\omega)}\right)^{1/2}\;\;\;\text{at $2\omega$}.$
@@
However, this assumption is unreliable since $\eta^{(2\omega)}_{1}$, $\eta^{(2\omega)}_{2}$, $\eta^{(\omega)}_{1}$, and $\eta^{(\omega)}_{2}$ can all differ substantially. This “degeneracy” constraint introduces significant uncertainty and yields a biased estimator of on-chip performance. In particular, when $\mathcal{S}_{\text{ON}}$ or $\mathcal{E}_{\text{ON}}$ is inferred by back-calculating from measured off-chip quantities $\mathcal{S}_{\text{OFF}}$ and $\mathcal{E}_{\text{OFF}}$ (Eq. (1)), the use of $x_{2}=\hat{\eta}_{2,\text{C}}$ under conventional calibration introduces systematic error.
@@
Figure 4: MC simulations of squeezing and SHG efficiency. Histograms of (a) off-chip squeezing and (b) SHG efficiency. The estimated (c) on-chip squeezing and (d) SHG efficiency. Black dashed lines mark the ground-truth of on-chip squeezing and SHG efficiency: $\mathcal{S}_{\text{ON}}(\eta^{(2\omega)}_{1})$ and $\mathcal{E}_{\text{ON}}(\eta^{(\omega)}_{2})$. In (c) and (d), the brown histograms represent conventional calibration assuming “degenerate” coupling efficiencies $\hat{\eta}_{1,C}=\hat{\eta}_{2,C}=0.784$ (pumping at $\omega$), while the green histograms show the estimates of BNOT.
@@
where $e^{2}_{\text{sqz}}(x_{1},x_{2})\equiv\left|1-\mathcal{S}_{\text{OFF}}(x_{1},x_{2})/\mathcal{S}_{\text{OFF}}(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})\right|^{2}$ and $e^{2}_{\text{shg}}(x_{1},x_{2})\equiv\left|1-\mathcal{E}_{\text{OFF}}(x_{1},x_{2})/\mathcal{E}_{\text{OFF}}(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})\right|^{2}$. In words, the coupling efficiencies $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ are estimated by determining the pair of values $\hat{\eta}_{1,\text{B}},\hat{\eta}_{2,\text{B}}$ that minimizes the equally weighted mean square error rate for the squeezing and SHG efficiency estimation. To evaluate the estimation outcome with BNOT, we consider the ground-truth interface efficiencies:
@@
[…]
and adopt the Monte Carlo (MC) simulation, using the physical parameters of the PPLN waveguide in Tab. 1. Fig. 2(c) shows histograms of $\hat{\eta}_{1,\text{B}}$ and $\hat{\eta}_{2,\text{B}}$ obtained from MC simulations. The close agreement with the ground-truth values confirms the unbiasedness of our estimators. However, with the conventional calibration approach, the estimation is $\hat{\eta}_{1,\text{C}}=\hat{\eta}_{2,\text{C}}=0.784$ for pump at $\omega$ or $0.741$ for pump at $2\omega$, both of which deviate from the pair of ground-truth efficiencies: $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ (black dashed vertical lines in Fig. 2(c)) that are relevant to squeezing and SHG experiments.
@@
To assess general ground-truth efficiencies, we scanned the full range of $\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2}\in\left[0,1\right]$. For each pair $(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})$, our estimator is applied using MC simulations, with the total root-mean-square error (RMSE) of $\hat{\eta}_{1,\text{B}}$ and $\hat{\eta}_{2,\text{B}}$, $\sqrt{e_{1}^{2}+e_{2}^{2}}$, shown in Fig. 3(a). Here, $e_{1}\equiv\sqrt{\mathbb{E}\left[(\hat{\eta}_{1,\text{B}}-\eta_{1}^{(2\omega)})^{2}\right]}$ and $e_{2}\equiv\sqrt{\mathbb{E}\left[(\hat{\eta}_{2,\text{B}}-\eta_{2}^{(\omega)})^{2}\right]}$ denote the RMSEs of $\hat{\eta}_{1,\text{B}}$ and $\hat{\eta}_{2,\text{B}}$, and $\mathbb{E}[\cdot]$ denotes the arithmetic mean in simulation trials. Five representative pairs of $(\eta_{1}^{(2\omega)},\eta_{2}^{(\omega)})$, indicated by the colored stars in Fig. 3(a). The corresponding RMSEs, $e_{1}$ and $e_{2}$, are shown in Fig. 3(b,c).
From Fig. 3(a), achieving a low RMSE imposes a much stricter requirement on the estimation of $\eta_{2}^{(\omega)}$ than on $\eta_{1}^{(2\omega)}$. For instance, the total estimation error remains $\leq 0.01$ (red dashed curve) only when $\eta_{1}^{(2\omega)}\geq 0.3$ and $\eta_{2}^{(\omega)}\geq 0.85$. This asymmetry arises because the detected squeezing depends nonlinearly on the output efficiency $\eta_{2}^{(\omega)}$, which enters exponentially in the measured quadrature variance. As a result, small deviations in $\eta_{2}^{(\omega)}$ strongly affect the observed squeezing. In contrast, SHG depends linearly on $\eta_{2}^{(\omega)}$ and quadratically on $\eta_{1}^{(2\omega)}$, distributing the effect of measurement noise $\xi_{\text{shg}}$ more symmetrically across both interfaces.
@@
Using the ground-truth efficiencies defined in Eq. (5), we estimate $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ with both the conventional calibration and BNOT methods. These values are then used to back-calculate the on-chip squeezing and SHG efficiency via the noiseless relations in Eq. (1). Based on the simulated off-chip data in Fig. 4(a,b), the reconstructed on-chip performances are shown in Fig. 4(c,d) with the associated ground truth (black dashed vertical lines). As expected, the BNOT estimator remains unbiased, whereas the conventional calibration exhibits systematic bias. Because BNOT jointly fits the nonlinear processes, the estimation of $\eta_{2}^{(\omega)}$, which strongly affects off-chip squeezing, directly constrains $\mathcal{S}_{\text{ON}}$ through a shared likelihood function. This self-consistency reduces both bias and variance in the reconstructed on-chip squeezing and SHG efficiency. Conversely, the conventional approach treats the two calibration steps independently, leading to systematically decorrelated estimates. This is most pivotal in high-squeezing regimes (e.g., $\mathcal{S}_{\text{OFF}}\sim 15\penalty 10000\ \text{dB}$), where BNOT’s unbiased posteriors for $(\eta_{1}^{(2\omega)},\eta_{2}^{(\omega)})$ translate directly into precise interface-efficiency requirements and the corresponding $\mathcal{S}_{\text{ON}}$ needed to reach the target (see Appendix).
II Outlook

### v2

@@ Abstract
Accurate evaluation of nonlinear photonic integrated circuits requires separating input and output coupling efficiencies (i.e., $\eta_{1}$ and $\eta_{2}$), yet the conventional linear-transmission calibration method recovers only their product (i.e., $\eta_{1}\,\eta_{2}$) and therefore introduces systematic bias when inferring on-chip performance from off-chip data. We present bidirectional nonlinear optical tomography (BNOT), a direction-aware metrology that uses forward and backward pumping of complementary nonlinear probes, with process-appropriate detection, to break the “degeneracy” of $\eta_{1}\,\eta_{2}$ and estimate individual interface efficiencies with tight confidence intervals. The method links off-chip measurements to on-chip quantities through a compact observation model that explicitly incorporates pump fluctuations and detector noise, and it frames efficiency extraction as a joint constrained optimization. Monte Carlo studies show unbiased convergence of the estimated efficiencies to ground truth with low error across realistic operating regimes. Using these efficiency estimates to reconstruct on-chip nonlinear figures of merit yields distributions centered on the true values with reduced variance, whereas conventional “degenerate” calibration is biased and can substantially misestimate on-chip performance. BNOT is hardware-compatible and platform-agnostic, and provides unbiased characterization of off- and on-chip coupling efficiencies across nonlinear processes, enabling reproducible, coupling-resolved benchmarking for scalable systems in quantum optics, frequency conversion, and precision metrology.
@@ I Introduction
Nonlinear photonic integrated circuits (PICs) promise a compact wafer-scale platform that transforms passive chips into active nonlinear engines for frequency conversion and generation of quantum states [1, 2]. Across emerging platforms such as lithium niobate (LiN), silicon nitride (SiN), aluminum nitride, and gallium arsenide, canonical processes, including spontaneous parametric down conversion (SPDC) [3, 4, 5], four-wave mixing (FWM) [6, 7, 8, 9, 10, 11, 12, 13], harmonic generation [14, 15, 16, 17], self-phase modulation (SPM) [18], optical parametric oscillation (OPO) [19, 20, 21, 22] and optical parametric amplification (OPA) [23, 3, 16], are now accessible on chip, enabling applications in spectroscopy, sensing, and quantum information science [3, 6, 7, 8, 9, 14, 15, 19, 20, 24]. These processes are carried out in diverse PIC structures (Fig. 1(a)), ranging from waveguides to a variety of microring resonators, which provide tailored dispersion and mode confinement. Multimode and microcomb architectures further scale brightness and mode count with favorable hardware efficiency [25], positioning nonlinear PICs as strong candidates for continuous-variable (CV) quantum computing [7, 26, 27] and classical optical computing [28, 29, 30, 31].
@@
A major challenge in nonlinear PICs is the lack of bias-free performance estimation, especially as nonlinear interaction strength increases. Because experiments typically rely on off-chip lasers and detectors, the reported metrics involve two coupling interfaces whose efficiencies $(\eta_{1},\eta_{2})$ are often unequal due to mode mismatch, polarization sensitivity, fabrication imperfections, or differing coupling schemes (e.g., edge coupler on one side and grating coupler on the other). Conventional linear-transmission measurements yield only the product $\propto\eta_{1}\,\eta_{2}$, obscuring which interface dominates the loss and leading to the “degenerate” estimate $(\eta_{1}\,\eta_{2})^{1/2}$, a source of systematic bias in evaluating on-chip performance [32, 33, 34, 32]. An example is chip-based second harmonic generation (SHG) measured off chip, where the detected efficiency depends asymmetrically on the two interfaces. This can be expressed as $f_{AS}(\eta_{1},\eta_{2})$, satisfying $f_{AS}(\eta_{1},\eta_{2})\neq f_{AS}(\eta_{2},\eta_{1})$ and scaling as $P_{\text{SHG}}/P_{\text{in}}^{2}\propto\eta_{\text{in}}^{2}\eta_{\text{out}}$ [35]. Conventional calibration assumes $\eta_{\text{in}}=\eta_{\text{out}}=(\eta_{1}\,\eta_{2})^{1/2}$, giving $P_{\text{SHG}}/P_{\text{in}}^{2}\propto(\eta_{1}\,\eta_{2})^{3/2}$ (curve (i) in Fig. 1(b)), while the actual response depends on direction: forward pumping ($\eta_{\text{in}}=\eta_{1}$, $\eta_{\text{out}}=\eta_{2}$) yields $\propto\eta_{1}^{2}\,\eta_{2}$ (curve (ii)), and backward pumping ($\eta_{\text{in}}=\eta_{2}$, $\eta_{\text{out}}=\eta_{1}$) yields $\propto\eta_{2}^{2}\,\eta_{1}$ (curve (iii)). This intrinsic asymmetry ultimately explains the systematic deviation observed under symmetric calibration. Wang et al. [36] used forward and backward SHG as symmetric and asymmetric references to attribute performance differences to the different input/output coupling efficiencies. However, such approaches still assume equal efficiencies in the symmetric case and cannot uniquely resolve individual efficiencies.
@@
Among nonlinear processes, squeezed light provides another stringent benchmark for on-chip device performance. Bulk-optics squeezing experiments have reached $15\penalty 10000\ \text{dB}$ [37], while envisioned applications require near-$10\penalty 10000\ \text{dB}$ for CV fault-tolerant quantum computing [26], gravitational-wave detectors [38], and Gottesman-Kitaev-Preskill (GKP) sources [39]. Integrated platforms have advanced from $1.7\penalty 10000\ \text{dB}$ demonstrations [10] to multi-decibel squeezing in LiN [7], cavity optomechanics [40], and foundry-compatible SiN [41, 12], with further gains in nanophotonic molecules and microcombs [42, 11]. The measured squeezing levels reported (off-chip) now exceed $4.9\penalty 10000\ \text{dB}$ in PPLN waveguides, $3.5\penalty 10000\ \text{dB}$ in Kerr microrings, $5.6\penalty 10000\ \text{dB}$ in microcombs [3, 6, 9], and attain $3.1\penalty 10000\ \text{dB}$ in wafer-scale integration [12]. As measured values increase and target chip squeezing thresholds approach $10-15\penalty 10000\ \text{dB}$, an accurate estimation of interface efficiencies $(\eta_{1},\eta_{2})$ becomes even more critical [43, 44, 6, 45].
@@
Figure 2: Histograms of Monte Carlo (MC) simulations with the ground-truth interface efficiencies in Eq. (4) and parameters in Tab. 1. (a) Simulated off-chip squeezing and (b) SHG efficiency. (c) BNOT estimates of $\hat{\eta}_{1,\,\text{B}}$ (blue) and $\hat{\eta}_{2,\,\text{B}}$ (red), with shaded regions showing 95 % confidence intervals (2.5–97.5 percentiles).
II Concept
Our BNOT methodology leverages the reversibility of two nonlinear processes within the same photonic device by pumping it in opposite directions. To benchmark its performance, we consider on-chip squeezed-light generation and SHG in an exemplary platform: periodically poled LiN (PPLN) waveguides. The calibration scheme is illustrated in Fig. 1(c). A pump field at $2\omega$ is injected into the PPLN waveguide through the left interface in the forward direction (left-to-right), while a pump at $\omega$ is simultaneously launched through the right interface in the backward direction (right-to-left). The corresponding interface coupling efficiencies are denoted $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$, where the superscript denotes the pump frequency. This bidirectional pumping exploits the reversibility of the nonlinear interactions, enabling both degenerate SPDC and SHG to occur concurrently. The corresponding output fields are measured at their respective ports using balanced homodyne detection and direct detection.
To relate on- and off-chip performances in squeezing and SHG, we define two estimation variables, $x_{1}$ and $x_{2}$, corresponding to the left- and right-coupling efficiencies of the PPLN waveguide. In particular, these are not the ground-truth efficiencies: $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ but the tunable parameters in our model used to recover them. The off-chip squeezing level (in dB) and SHG efficiency (in W-1m-2) are expressed as
@@ II Concept
$\mathcal{E}_{\text{OFF}}\,(x_{1},x_{2})$ $=x_{1}\,\mathcal{E}^{(\varepsilon_{2}\,,\,\Delta k)}_{\text{ON}}(x_{2})+\xi_{\text{shg}},$
where $\mathcal{S}^{(\varepsilon_{1},\,\Delta k)}_{\text{ON}}$ and $\mathcal{E}^{(\varepsilon_{2},\,\Delta k)}_{\text{ON}}$ denote the analytical models for the on-chip squeezing [46] and SHG efficiency [17], respectively, with their derivations provided in App. A. The parameters $\varepsilon_{1}\sim U(1-\delta_{1},1+\delta_{1})$ and $\varepsilon_{2}\sim U(1-\delta_{2},1+\delta_{2})$ describe the uniform fractional fluctuations in the pump powers, where $\delta_{1}$ and $\delta_{2}$ specify the corresponding half-widths of the fluctuation ranges. The wave-vector mismatch is modeled as $\Delta k\sim\mathcal{N}(0,\sigma_{\Delta k}^{2})$, with $\sigma_{\Delta k}$ representing the standard deviation arising from the temperature instability of the PPLN waveguide. Additional measurement noise in the squeezing and SHG signals is captured by $\xi_{\text{sqz}}\sim\mathcal{N}(0,\sigma_{\text{sqz}}^{2})$ and $\xi_{\text{shg}}\sim\mathcal{N}(0,\sigma_{\text{shg}}^{2})$, which account for residual technical fluctuations in the detection system.
@@
Figure 3: (a) Total mean-square error (MSE) map of efficiency estimation using BNOT, $e_{1}^{2}+e_{2}^{2}$, where $e^{2}_{1}$ and $e^{2}_{2}$ are the estimator’s MSEs of $\eta^{(2\omega)}_{1}$ and $\eta^{(\omega)}_{2}$, respectively. The red dashed contour marks the 20-20 dB boundary. Colored stars indicate the parameter sets used for Monte Carlo simulations. (b) MSE of $\hat{\eta}_{1,\,\text{B}}$ and (c) MSE of $\hat{\eta}_{2,\,\text{B}}$ as functions of the number of independent trials, with colors matching the parameter sets in (a).
@@
The conventional linear transmission-based calibration approach cannot distinguish between left-to-right and right-to-left pumping [32, 33, 34, 32]. As a result, it enforces a “degenerate”-interface assumption, estimating the efficiencies as
$\hat{\eta}_{1,\text{C}}=\hat{\eta}_{2,\text{C}}=\mathcal{N}\left(\sqrt{\eta_{1}^{(\omega)}\,\eta_{2}^{(\omega)}},\sigma_{\text{C}}^{2}\right),$ (2)
@@
with the measurement standard deviation $\sigma_{\text{C}}$. Clearly, the mean of the estimate is biased by the ground-truth efficiencies that are relevant to squeezing and SHG experiments (i.e., $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$). This “degeneracy” constraint introduces significant uncertainty and yields a biased estimator of on-chip performance. In particular, when the on-chip $\mathcal{S}_{\text{ON}}$ or $\mathcal{E}_{\text{ON}}$ is inferred by back-calculating the measured quantities off the chip $\mathcal{S}_{\text{OFF}}$ and $\mathcal{E}_{\text{OFF}}$ (Eq. (1)), the conventional calibration procedure replaces $x_{2}$ with the estimated coupling efficiency $\hat{\eta}_{2,\text{C}}$, thus introducing a systematic bias in the inferred result.
@@
Figure 4: MC simulations of squeezing and SHG efficiency. Histograms of (a) off-chip squeezing and (b) SHG efficiency. The estimated (c) on-chip squeezing and (d) SHG efficiency are shown. Black dashed lines mark the ground-truth of on-chip squeezing and SHG efficiency: $\mathcal{S}^{(\varepsilon_{1},\,\Delta k)}_{\text{ON}}$ and $\mathcal{E}^{(\varepsilon_{2},\,\Delta k)}_{\text{ON}}$. In (c) and (d), the brown histograms represent conventional calibration assuming “degenerate” coupling efficiencies, while the green histograms show the estimates of BNOT.
III Simulation Results
@@
where $e^{2}_{\text{sqz}}(x_{1},x_{2})\equiv\left|1-\mathcal{S}_{\text{OFF}}(x_{1},x_{2})/\mathcal{S}_{\text{OFF}}(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})\right|^{2}$ and $e^{2}_{\text{shg}}(x_{1},x_{2})\equiv\left|1-\mathcal{E}_{\text{OFF}}(x_{1},x_{2})/\mathcal{E}_{\text{OFF}}(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})\right|^{2}$. In words, the coupling efficiencies $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ are estimated by determining the pair of values $\hat{\eta}_{1,\,\text{B}},\hat{\eta}_{2,\,\text{B}}$ that minimizes the equally weighted mean square error rate for the squeezing and SHG efficiency estimation.
We consider the ground-truth interface efficiencies:
@@
and adopt the Monte Carlo (MC) simulation in Fig. 2(a, b), using the physical parameters of the PPLN waveguide in Tab. 1. Fig. 2(c) presents the histograms of the estimators $\hat{\eta}_{1,\,\text{B}}$ and $\hat{\eta}_{2,\,\text{B}}$ obtained from the MC simulations. Their close agreement with the ground-truth values confirms the unbiasedness of the proposed estimators. Moreover, Fig. 2(c) reveals that the estimation of $\eta_{2}^{(\omega)}$ exhibits a significantly narrower confidence interval than that of $\eta_{1}^{(2\omega)}$. This asymmetry arises because, in squeezing, the output efficiency $\eta_{2}^{(\omega)}$ directly sets the amount of vacuum noise mixed into the detected field, so small changes in $\eta_{2}^{(\omega)}$ cause large variations in the observed squeezing. In contrast, the SHG signal depends linearly on $\eta_{2}^{(\omega)}$ and quadratically on $\eta_{1}^{(2\omega)}$, which distributes the measurement noise $\xi_{\text{shg}}$ more evenly between the two interfaces.
@@
To assess general ground-truth efficiencies, we scanned the full range of $\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2}\in\left[0,1\right]$. For each pair $(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})$, our estimator is applied using MC simulations, with the overall mean-square error (MSE) of $\hat{\eta}_{1,\,\text{B}}$ and $\hat{\eta}_{2,\,\text{B}}$, $e_{1}^{2}+e_{2}^{2}$, in decibels, shown in Fig. 3(a) (the red dashed curve marks the MSE boundary of 20-20 dB). Here, $e_{1}\equiv\sqrt{\mathbb{E}\left[(\hat{\eta}_{1,\,\text{B}}-\eta_{1}^{(2\omega)})^{2}\right]}$ and $e_{2}\equiv\sqrt{\mathbb{E}\left[(\hat{\eta}_{2,\,\text{B}}-\eta_{2}^{(\omega)})^{2}\right]}$, where $\mathbb{E}[\cdot]$ denotes the arithmetic mean in simulation trials. Five representative pairs of $(\eta_{1}^{(2\omega)},\eta_{2}^{(\omega)})$, indicated by the colored stars in Fig. 3(a). The corresponding MSEs, $e^{2}_{1}$ and $e^{2}_{2}$, are shown in Fig. 3(b,c).
@@
Using the ground-truth efficiencies defined in Eq. (4), we estimate $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ with both conventional calibration and BNOT methods. These values are then used to back-calculate the on-chip squeezing and SHG efficiency via the noiseless relations in Eq. (1). Based on the simulated off-chip data in Fig. 4(a,b), the reconstructed on-chip performances are shown in Fig. 4(c,d) with the associated ground truth (black dashed vertical lines). From the simulation result, the BNOT estimator remains unbiased, whereas the conventional calibration exhibits a systematic bias. Because BNOT jointly fits the nonlinear processes, the estimation of $\eta_{2}^{(\omega)}$, which strongly affects off-chip squeezing, directly constrains $\mathcal{S}_{\text{ON}}$ through a shared likelihood function. This self-consistency reduces both bias and variance in the reconstructed on-chip squeezing and SHG efficiency. In contrast, the conventional approach treats the two calibration steps independently, leading to systematically decorrelated estimates. This is most pivotal in high-squeezing regimes (e.g., $\mathcal{S}_{\text{OFF}}\sim 15\penalty 10000\ \text{dB}$), where BNOT’s unbiased posteriors for $(\eta_{1}^{(2\omega)},\eta_{2}^{(\omega)})$ translate directly into precise interface-efficiency requirements and the corresponding $\mathcal{S}_{\text{ON}}$ needed to reach the target (see App. B).
IV Conclusion and outlook